OpenStudy (anonymous):

–4x + 3y = 6. help me graph this plz!!

OpenStudy (anonymous):

Easiest way to graph a linear equation is slope-intercept form. @Nana3456, do you know how to write this in slope-int form?

OpenStudy (anonymous):

the graph is a line joining points(-3/2,0) and (0,2)

OpenStudy (anonymous):

if someone could draw like ta little graph so I can see plz??

OpenStudy (anonymous):

-4x + 3y = 6 3y = 4x + 6 y = 4/3 x + 2 So, the y-intercept is (0,2) and the slope is 4/3. Now, can you graph this?

OpenStudy (anonymous):

solve for y and just copy it into google

OpenStudy (anonymous):

4/3 x + 2 that to google

OpenStudy (anonymous):